The parallel rails with negligible are 14.5 cm apa
Answer #1

flux 1 = B*A1


A1 = area = l*dx


l = 0.145 m

induced emf 1 = B*A1/dt = B*l*dx/dt = B*l*v1 = 0.01*0.145*4 = 0.0058 v


flux 2 = B*A2


A2 = area = l*dx


l = 0.145 m

induced emf 2 = B*A2/dt = B*l*dx/dt = B*l*v2 = 0.01*0.145*2 = 0.0029 v

for left loop

E1 - I1*R1 - I3*R3 = 0

0.0058 - I1*13.5 - I3*5 = 0.........(1)


for right loop


E2 - I2*R2 + I3*R3 = 0


0.0029 - (I2*15) + (I3*5) = 0.......(2)

I1 = I2 + I3 ....(3)


I1 = 378 uA


I2 = 240 uA

I3 = 138 uA <<<<_---------answer